OpenStudy (anonymous):

Will be a fan if you can help me with this i am in the 6th grade so not so hard. ] Based on simple interest, find the total amount saved or owed. principal = $1700; rate = 6%; time = 5 years A. $1802 B. $1904 C. $2210 D. $2312

OpenStudy (nurali):

I = PRT I = 1700 * 0.06 * 5 I = 510 1700 + 510 = ?

OpenStudy (anonymous):

2,210 Thank you so much kind sir like i said ill fan you!